OpenStudy (anonymous):

Factor completely: 18x^2 - 21x - 15 A. 3(2x + 1)(3x - 5) B. 3(2x - 5)(3x + 1) C. 3(2x - 1)(3x + 5) D. 3(6x + 1)(x - 5)

OpenStudy (anonymous):

the middle term could be different A) 3 (-10x +3x) = 3 (-7x) = -21x B. 3 (2x+-15x) = 3(-13x) = -26x and so on...

OpenStudy (anonymous):

C) 3(10x-3x ) = 3 (7x) = 21x D) 3(-30x+x) = 3(-29x) = too high

OpenStudy (anonymous):